# Text-Tuple-Table: Towards Information Integration in Text-to-Table Generation via Global Tuple Extraction

### v1

@@ C.2 Zero-Shot Setting
The zero-shot and chain-of-thought inference for LLaMA-2 models are conducted on eight V100 GPUs. When obtaining outputs from LLMs via APIs, we ensure deterministic results by setting the temperature to zero. In Table 5, we list the model names corresponding to different model types. Note that unless specifically stated otherwise, when a model type from the table is used, the model name corresponding to it in the table is the one being invoked.
@@
Table 5: Model types and their corresponding names of LLMs used in our experiments.
C.3 Prompts
@@ F.3 Results Analysis
The results of this experiment are presented in Figure 5. As discussed in Section 6.3.2, T3 demonstrates higher Auto-QA Coverage than the baseline, proving its ability to generate higher quality tables in text-to-table generation tasks without instructions. We also experiment with using the tuples extracted in the first step of T3 directly as the generated tables and assessed their coverage, resulting in constructive findings. The coverage curve of T2 is entirely above that of T3, indicating a certain loss of information from tuple to table. Although this could also be due to insufficient table question answering capabilities leading to a decrease in metrics, it is necessary to employ relevant techniques Wu et al. (2023); Wang et al. (2024) to mitigate the impact of this factor in future work. Nevertheless, the conclusion that T3 improves the performance of the previous work remains sound.

### v2

@@ C.2 Zero-Shot Setting
The zero-shot and chain-of-thought inference for LLaMA-2 models are conducted on eight V100 GPUs Wang et al. (2024a). When obtaining outputs from LLMs via APIs, we ensure deterministic results by setting the temperature to zero. In Table 6, we list the model names corresponding to different model types. Note that unless specifically stated otherwise, when a model type from the table is used, the model name corresponding to it in the table is the one being invoked.
@@
Table 6: Model types and their corresponding names of LLMs used in our experiments under zero-shot setting.
C.3 Few-Shot Setting
Because the input lengths of LiveSum dataset generally exceed 2,000 tokens, even testing with 1-shot would surpass the maximum token limit (4,096) of most LLMs. Therefore, we select two models, ChatGPT-16k OpenAI (2022) and GPT-4o OpenAI (2024b), for supplementary experiments Regarding the selection of few-shot examples, we randomly choose samples from the training set where all events occur. In the Chain of Thought setting, the reasoning process is included in the examples.
Model Type Model Name
ChatGPT-16k gpt-3.5-turbo-16k-0613
GPT-4o gpt-4o-2024-05-13
Table 7: Model types and their corresponding names of LLMs used in our experiments under few-shot setting.
C.4 Prompts
@@ F.3 Results Analysis
The results of this experiment are presented in Figure 5. As discussed in Section 6.3.2, T3 demonstrates higher Auto-QA Coverage than the baseline, proving its ability to generate higher quality tables in text-to-table generation tasks without instructions. We also experiment with using the tuples extracted in the first step of T3 directly as the generated tables and assessed their coverage, resulting in constructive findings. The coverage curve of T2 is entirely above that of T3, indicating a certain loss of information from tuple to table. Although this could also be due to insufficient table question answering capabilities leading to a decrease in metrics, it is necessary to employ relevant techniques Wu et al. (2023); Wang et al. (2024d) to mitigate the impact of this factor in future work. Nevertheless, the conclusion that T3 improves the performance of the previous work remains sound.
F.4 Error Analysis
We carry out a supplementary case study on the Wiki40B dataset. We sample 50 articles and observe the following situations:
  • In the first phase, the average accuracy of the tuples obtained from the text-to-tuple process is 81.6%. We define erroneous tuples as those containing: (i) direct extractions of original sentences, (ii) incomplete triples, and (iii) duplicate triples.
  • In the second phase, tuple integration, 24% of the articles undergo integration of similar items, while the remainder simply involves sorting and categorizing the tuples or outputting them unchanged.
  • In the third phase, tuple-to-table essentially involves treating each tuple as a row to be filled into a table, but two scenarios may introduce errors: (i) 14% of the articles involve the fragmentation of long sentences, leaving only incomplete tuples, and (ii) 6% of the articles continue to merge similar items based on the tuples, combining related tuples into long sentences. (While this step may not necessarily introduce errors, it could increase the difficulty of reasoning, and the AutoQA metric may decrease.)